# A variational formulation of the free energy of mixed quantum-classical systems: coupling classical and electronic density functional theories

### v5

@@ II.2 Mixed quantum-classical equilibrium density, partition function and Helmholtz free energy
From a more formal point of view, it is worth noticing that Eq. (9) corresponds to the real-space representation of the EDM, originally obtained by Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001) in the adiabatic basis, i.e. in the basis of eigenvectors of $\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$. In this basis, the density matrix is diagonal , as found by Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001). In the case of quantum electrons and classical nuclei, the adiabatic states are the eigenstates of the Born-Oppenheimer Hamiltonian. Therefore, Eq. (9) ignores the non-adiabatic couplings between these states. In the context of ab initio MD, this implies that $\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})$ can be obtained from classical nuclear trajectories, with forces averaged over all adiabatic electronic states weighted by the corresponding Boltzmann factors. Consequently, it entirely ignores surface hopping between different adiabatic states, which is why $\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})$ is referred to as the adiabatic QM/MM density matrix. There is a flourishing literature addressing methods going beyond this adiabatic approximation, specially in the context of quantum dynamics or statistical mechanics for QM/MM systems (see for instance Ref. Kapral, 2015 for a review), but treatment of non adiabatic couplings is beyond the scope of this paper.
@@
It is noteworthy that, for each value of $\bf{Q}$}, $\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}({\bf{Q}})$ defined in Eq. (23) and (24) is a density matrix, i.e. positive definite and of unity norm. Similarly, $\boldsymbol{p}_{\text{eq}}$ of Eq. (26) is a probability distribution. Therefore, $\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}$ and $\boldsymbol{p}_{\text{eq}}$ are special cases of more general objects: density matrices representing $N_{{\text{qm}}}$ quantum particles and probability distribution functions representing $N_{{\text{mm}}}$ classical particles, respectively. This will allow to formulate a variational formulation of the Helmholtz free energy for a QM/MM system in Sec. II.3.
@@
II.3 Variational formulation of the Helmholtz free energy for a QM/MM system
@@ II.3.1 Variational formulation
II.3.2 Rewriting of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$
@@
$\quad\int\text{d}{\bf{q}}\text{d}{\bf{q}}^{\prime}(\bf{t}({\bf{q}},{\bf{q}}^{\prime})+{w}_{{\text{qm}}}({\bf{q}},{\bf{q}}^{\prime})){\boldsymbol{\rho}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}}),$
@@ III.2 Explicit coupling in terms of eDFT and cDFT
where we can explicitly recognize the purely quantum and purely classical free-energy functionals, together with the QM/MM coupling expressed as the sum of the mean-field and correlation terms. If $\delta\mathcal{F}_{{\text{qm}}}$ is ignored, we recover the mean-field coupling between the eDFT and cDFT functionals that is generally used in the literaturePetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Tang et al. (2020); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024). We highlight here that, although the outcome of the present derivation is rather natural as it consists of a multi component DFT formalism, the explicit definition of $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ in Eq. (64) can help in future developments of approximations for $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$. For instance, the variational formulation in terms of full-body densities allows the use the adiabatic connection to reformulate the problem, and therefore potentially to import several successful strategies used in eDFT to approximate correlation functionals.
@@ III.3 Applications to solvation problems
In Petrosyan et al’s joint DFT Petrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007), the interaction of the solvent with both the solute nuclei and electrons is dealt as a whole and the coupling is described by the use of pseudo potential. Thus Eq. (67) solely contains the quantum functional $\mathcal{F}_{\text{qm}}$, the solvent one $\mathcal{F}_{\text{mm}}$ and $\varepsilon_{{\text{qm}}}^{{\text{mm}}}$. The coupling term is modeled as a dielectric screening due to a non local dielectric function depending on the electron density and a mean-field term between both densities interacting through a Gaussian repulsive potential.
@@
In our previous worksJeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024), we also ignored the correlation functional $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}$ in Eq. (65) and limit ourselves to the mean-field term. The interaction potential was also the sum of a Lennard-Jones and a Coulombic term. However, we do not resort to point charges representation to compute the electrostatic energy but directly use the electronic density
@@
We also proposed to couple wave function theory-based methods, such as Hartree-Fock and selected configuration interaction with cDFTLabat, Giner, and Jeanmairet (2024) to approximate the universal Levy-Lieb functional $\mathcal{F}_{\text{qm}}[\rho]$ of Eq. (65).
The results obtained in these worksPetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Tang et al. (2020); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024) clearly indicate that, for ground state properties such as dipoles for instance, the mean-field QM/MM approximation gives already a decent improvement with respect to in vaccuum pure QM calculations.
@@ IV Conclusion
The assumptions of these derivations is a QM/MM partitioning of the system, together with models for both the MM part and the QM/MM interaction. We therefore do not discuss the QM/MM crossover nor how to go beyond the classical models. The starting point of these derivations is the proposal of the QM/MM equivalent to the quantum ground state variational principle or the classical Gibbs measure variational principle (see Sec. II.3). This QM/MM functional is formalized in terms of a $N_{{\text{mm}}}$ classical probability density and an $N_{{\text{qm}}}$ quantum density matrix which also parametrically depends on the $N_{{\text{mm}}}$ classical variables (see Sec. Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ for proofs of the variational property). The latter was proposed thanks to the knowledge of the equilibrium density of the QM/MM system (see Eqs. (14) and (17)) which can be obtained from first principle through a semi classical expansion of a partial Wigner transformation (see Ref. Nielsen, Kapral, and Ciccotti, 2001 and the section II of the supplementary information for a detailed derivation).
Starting from this QM/MM variational principle written in terms of $N$-body objects and thanks to constrained search formalisms similar to that proposed by Levy and Lieb in eDFT, we establish a universal QM/MM functional depending only on the one-body classical and quantum objects (see Sec. III.1). Building on these results, we explicitly make a coupling with the finite temperature eDFT and cDFT functionals, which allows us to rigorously link the two theories for the description of a QM/MM system (see Sec. III.2). We then propose a practical application of this theory in the case of solvation problems, and propose a mean-field treatment. Thanks to these derivations, it becomes clear that the explicit correlation between the classical and quantum electronic systems, were implicitly neglected in our previously reported QM/cDFT calculationsJeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024).

### v6

@@ II.3 Mixed quantum-classical equilibrium density, partition function and Helmholtz free energy
From a more formal point of view, it is worth noticing that Eq. (14) corresponds to the real-space representation of the EDM, originally obtained by Nielsen et. al.Nielsen et al. (2001) in the adiabatic basis, i.e. in the basis of eigenvectors of $\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$. In this basis, the density matrix is diagonal , as found by Nielsen et. al.Nielsen et al. (2001). In the case of quantum electrons and classical nuclei, the adiabatic states are the eigenstates of the Born-Oppenheimer Hamiltonian. Therefore, Eq. (14) ignores the non-adiabatic couplings between these states. In the context of ab initio MD, this implies that $\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})$ can be obtained from classical nuclear trajectories, with forces averaged over all adiabatic electronic states weighted by the corresponding Boltzmann factors. Consequently, it entirely ignores surface hopping between different adiabatic states, which is why $\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})$ is referred to as the adiabatic QM/MM density matrix. There is a flourishing literature addressing methods going beyond this adiabatic approximation, specially in the context of quantum dynamics or statistical mechanics for QM/MM systems (see for instance Ref. Kapral, 2015 for a review), but treatment of non adiabatic couplings is beyond the scope of this paper.
@@
It is noteworthy that, for each value of $\bf{Q}$}, $\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}({\bf{Q}})$ defined in Eq. (28) and (29) is a density matrix, i.e. positive definite and of unity norm. Similarly, $\boldsymbol{p}_{\text{eq}}$ of Eq. (31) is a probability distribution. Therefore, $\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}$ and $\boldsymbol{p}_{\text{eq}}$ are special cases of more general objects: density matrices representing $N_{{\text{qm}}}$ quantum particles and probability distribution functions representing $N_{{\text{mm}}}$ classical particles, respectively. This will allow to formulate a variational formulation of the Helmholtz free energy for a QM/MM system in Sec. II.4.
@@
II.4 Variational formulation of the Helmholtz free energy for a QM/MM system
@@ II.4.1 Variational formulation
II.4.2 Rewriting of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$
@@
$\quad\int\text{d}{\bf{q}}\text{d}{\bf{q}}^{\prime}(\bf{t}({\bf{q}},{\bf{q}}^{\prime})+\delta({\bf{q}}-{\bf{q}}^{\prime}){w}_{{\text{qm}}}({\bf{q}},{\bf{q}}^{\prime})){\boldsymbol{\rho}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}}),$
@@ III.2 Explicit coupling in terms of eDFT and cDFT
where we can explicitly recognize the purely quantum and purely classical free-energy functionals, together with the QM/MM coupling expressed as the sum of the mean-field and correlation terms. If $\delta\mathcal{F}_{{\text{qm}}}$ is ignored, we recover the mean-field coupling between the eDFT and cDFT functionals that is generally used in the literaturePetrosyan et al. (2005, 2007); Tang et al. (2020); Jeanmairet et al. (2020); Labat et al. (2024). We highlight here that, although the outcome of the present derivation is rather natural as it consists of a multi component DFT formalism, the explicit definition of $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ in Eq. (69) can help in future developments of approximations for $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$. For instance, the variational formulation in terms of full-body densities allows the use the adiabatic connection to reformulate the problem, and therefore potentially to import several successful strategies used in eDFT to approximate correlation functionals.
@@ III.3 Applications to solvation problems
In Petrosyan et al’s joint DFT Petrosyan et al. (2005, 2007), the interaction of the solvent with both the solute nuclei and electrons is dealt as a whole and the coupling is described by the use of pseudo potential. Thus Eq. (72) solely contains the quantum functional $\mathcal{F}_{\text{qm}}$, the solvent one $\mathcal{F}_{\text{mm}}$ and $\varepsilon_{{\text{qm}}}^{{\text{mm}}}$. The coupling term is modeled as a dielectric screening due to a non local dielectric function depending on the electron density and a mean-field term between both densities interacting through a Gaussian repulsive potential.
@@
In our previous worksJeanmairet et al. (2020); Labat et al. (2024), we also ignored the correlation functional $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}$ in Eq. (70) and limit ourselves to the mean-field term. The interaction potential was also the sum of a Lennard-Jones and a Coulombic term. However, we do not resort to point charges representation to compute the electrostatic energy but directly use the electronic density
@@
We also proposed to couple wave function theory-based methods, such as Hartree-Fock and selected configuration interaction with cDFTLabat et al. (2024) to approximate the universal Levy-Lieb functional $\mathcal{F}_{\text{qm}}[\rho]$ of Eq. (70).
The results obtained in these worksPetrosyan et al. (2005, 2007); Tang et al. (2020); Jeanmairet et al. (2020); Labat et al. (2024) clearly indicate that, for ground state properties such as dipoles for instance, the mean-field QM/MM approximation gives already a decent improvement with respect to in vaccuum pure QM calculations.
@@ IV Conclusion
The assumptions of these derivations is a QM/MM partitioning of the system, together with models for both the MM part and the QM/MM interaction. We therefore do not discuss the QM/MM crossover nor how to go beyond the classical models. The starting point of these derivations is the proposal of the QM/MM equivalent to the quantum ground state variational principle or the classical Gibbs measure variational principle (see Sec. II.4). This QM/MM functional is formalized in terms of a $N_{{\text{mm}}}$ classical probability density and an $N_{{\text{qm}}}$ quantum density matrix which also parametrically depends on the $N_{{\text{mm}}}$ classical variables (see Sec. Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ for proofs of the variational property). The latter was proposed thanks to the knowledge of the equilibrium density of the QM/MM system (see Eqs. (19) and (22)) which can be obtained from first principle through a semi classical expansion of a partial Wigner transformation (see Ref. Nielsen et al., 2001 and the section II of the supplementary information for a detailed derivation).
Starting from this QM/MM variational principle written in terms of $N$-body objects and thanks to constrained search formalisms similar to that proposed by Levy and Lieb in eDFT, we establish a universal QM/MM functional depending only on the one-body classical and quantum objects (see Sec. III.1). Building on these results, we explicitly make a coupling with the finite temperature eDFT and cDFT functionals, which allows us to rigorously link the two theories for the description of a QM/MM system (see Sec. III.2). We then propose a practical application of this theory in the case of solvation problems, and propose a mean-field treatment. Thanks to these derivations, it becomes clear that the explicit correlation between the classical and quantum electronic systems, were implicitly neglected in our previously reported QM/cDFT calculationsJeanmairet et al. (2020); Labat et al. (2024).